# Quantum property testing in sparse directed graphs

### v2

@@ 4.3 Main lower bound
Let us fix $f=(\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k})$ with domain $D=D_{(\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k})}$ (see Proposition 4.12). For technical reasons, in the rest of the section we fix $k\geq 3$ and $N=\lceil 20(2k)^{k/2}\rceil R$ [^2] These parameters are used in [polynomial_strikes_back] to prove Proposition A.4, which we will use.
@@
This subsection is dedicated to the proof of the following lemma which, together with Lemma 4.10, implies Theorem 4.1. Observe that we have to zero out the support of the dual polynomial outside of $H_{\leq N}^{NR}$, since our target domain is not $D$ but $D\cap H_{\leq N}^{NR}$ in Lemma 4.10.
@@
The construction of $\zeta$ starts by block composing (Definition 4.13) two dual polynomials $\phi,\psi$, one for $\mathrm{GapOR}_{R}^{\gamma}$ and one for $\mathrm{THR}_{N}^{k}$. The dual polynomial $\phi$ for $\mathrm{GapOR}_{R}^{\gamma}$ is given by Proposition 4.15. The dual polynomial $\psi$ for $\mathrm{THR}_{N}^{k}$ is given by the first part of Lemma 4.16.
The block composition $\phi\star\psi$ is a good candidate for the dual polynomial of $f$. Indeed, Lemma 4.18 shows that it satisfies Equation 8, showing correlation at least $9/10>2/3$. One could also check that it satisfies Equation 9. Nonetheless it does not satisfy Equation 7.
We can now use the second part of Lemma 4.16 to argue that there exists another dual polynomial $\zeta$ that satisfies Equation 7 and Equation 9. Moreover, this $\zeta$ is close to $\phi\star\psi$ so that it also satisfies Equation 8, with the weaker but sufficient correlation $9/10-2/9>2/3$. This concludes the proof. ∎
@@
The second lemma is the rephrasing of several scattered results in [polynomial_strikes_back] that we unify into one statement for more clarity. For the sake of completeness, we explain how to prove it in Section A.2.
@@
Let $N=\lceil 20(2k)^{k/2}\rceil R$ and $\phi:\{-1,1\}^{R}\to\mathbb{R}$ from Proposition 4.15. Then there exist $\psi:\{-1,1\}^{N}\to\mathbb{R}$ and $\zeta:\{-1,1\}^{NR}\to\mathbb{R}$ such that
@@
For the next lemma, we will use the following proposition, which was implicitly used in the proofs of [polynomial_strikes_back, Propositions 5.5 and 5.6], but not stated in this general form. We include its proof in Section A.3. By convention, we denote $D_{+1}=D_{+}$ and $D_{-1}=D_{-}$.
@@
Now we focus on the composed function $f$ (and the dual composition $\phi\star\psi$), the domain is not restricted to small Hamming weight inputs anymore.
@@
Let $N=\lceil 20(2k)^{k/2}\rceil R$ and $0<\gamma<1/4^{k-1}$. Functions $\phi$ from Proposition 4.15 and $\psi$ from Lemma 4.16 satisfy
@@
Proof of claim. Using Item 2 of Proposition 4.17, the left hand side can be written as
@@
Recall that $\phi(z)=0$ when $z$ is anything but $-1^{R}$ or $1^{R}$, so only two terms are left to study.
If $z=-1^{R}$, using Item 1 of Lemma 4.16 each $y_{i}$ is $-1$ with probability $\leq 1-1/4^{k-1}$ and 1 with probability $\geq 1/4^{k-1}$. If there is any $y_{i}=1$ then the value of the OR is still $-1$. The probability of this event is $\geq 1-(1-1/4^{k-1})^{R}\geq 1-e^{-\frac{R}{4^{k-1}}}$. So the expected value is $\leq(-1)(1-e^{-\frac{R}{4^{k-1}}})+e^{-\frac{R}{4^{k-1}}}=-1+2e^{-\frac{R}{4^{k-1}}}$. Since in this case $\phi(-1^{R})=-1/2$, the contribution to the sum is at most $1/2-e^{-\frac{R}{4^{k-1}}}$.
If $z=1^{R}$, then, using Item 1 of Lemma 4.16 again, each $y_{i}$ is $-1$ with probability $\leq 1/(48N)$. If any $y_{i}$ is $-1$ then the value of the OR becomes $-1$. The union bound tells us that the probability of this is $\leq R/(48N)$, so the expected value is at least $-R/(48N)+1-R/(48N)=1-R/(24N)$. Multiplied by $\phi(1^{R})=1/2$ the contribution is at least $1/2-R/(48N)$. Thus, the first term can be lower bounded by $1-e^{-\frac{R}{4^{k-1}}}-\frac{R}{48N}$. $\diamond$
@@
Proof of claim. By Item 1 of Proposition 4.17 with $S=\{-1,1\}^{NR}\setminus D$, the term can be written as follows,
@@
which, using that $|\phi(z)|=1/2$ when $z$ is $-1^{R}$ or $1^{R}$ and 00 otherwise, collapses to
$=\Pr_{x\sim\lambda^{\otimes R}}[x\notin D\mid(\dots,\mathrm{sgn}(\psi(x_{i})),z\dots)=-1^{R}]+\Pr_{x\sim\lambda^{\otimes R}}[x\notin D\mid(\dots,\mathrm{sgn}(\psi(x_{i})),\dots)=1^{R}].$
@@
Similarly, for the second term, if we sample $x_{i}$ from the conditional distribution $(\lambda|\mathrm{sgn}(\psi(x_{i}))=-1)$, then $\Pr[q_{i}=1]=\Pr[\mathrm{THR}_{N}^{k}(x_{i})=1|\mathrm{sgn}(\psi(x_{i}))=-1]=2\sum_{x_{i}\in D_{-}}|\psi(x_{i})|\leq 1-\frac{1}{4^{k-1}}$ (for the last step we used Item 1 of Lemma 4.16 again).
Then, similarly to the first term, we can upper bound the probability. Now in the last step, we use the Chernoff bound, which introduces the constraint $\gamma<\frac{1}{4^{k-1}}$.
@@
Finally, we can conclude the proof of Theorem 4.1.
Proof of Theorem 4.1.
By Lemma 4.14, there is a dual polynomial for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ of pure high degree $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ that is only supported on $H_{\leq N}^{NR}$. By Theorem 4.8 this means that the double-promise $\delta$-approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain restricted to $H_{\leq N}^{NR}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$. Using Lemma 4.10 with $c=\lceil 20(2k)^{k/2}\rceil$, we obtain that the bounded-error quantum query complexity of $\mathrm{Collision}_{N,N}^{k,\gamma^{\prime}}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ if $\gamma^{\prime}=\gamma/c<1/(4^{k-1}\lceil 20(2k)^{k/2}\rceil)$. This implies the same lower bound on testing $k$-collision-freeness with $\varepsilon=\gamma^{\prime}$ as $\mathrm{Collision}$ is just a more restricted version of the same problem. ∎
@@ 5 Testing $3$-colorability
Theorem 5.1 (Restatement of Theorem 1.4).
@@ 5.1 Classical lower bound approach for testing $3$-colorability
As we discussed in the overview, to prove the lower bound of $3$-colorability, the authors in [bogdanov2002lower] studied another problem called E$(3,c)$LIN-2, a problem related to deciding the satisfiability of a system of linear equations. Then the authors designed a reduction to $3$-colorability from E$(3,c)$LIN-2, which finally proves the linear query complexity lower bound for testing $3$-colorability. We will also follow a similar approach here. Let us first formally define the problem of E$(3,c)$LIN-2, where below $\mathbb{F}_{2}$ denotes the field with two elements.
Definition 5.3 (E$(3,c)$LIN-2).
@@
Using the existence of the matrix $A$ corresponding to Lemma 5.4, the authors in [bogdanov2002lower] used Yao’s minimax lower bound technique to prove a linear lower bound for testing E$(3,c)$LIN-2. More formally, they designed a pair of hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$, such that unless $\Omega(N)$ queries are performed, no algorithm can distinguish between them.
@@
There exists a matrix $A\in\{0,1\}^{cN\times N}$ (similar to the matrix mentioned in Lemma 5.4) such that given a parameter $\varepsilon\in(0,1)$ and query access to $A$ and a vector $y\in\{0,1\}^{cN}$, in order to distinguish if there exists another vector $x\in\{0,1\}^{N}$ such that $Ax=y$, or for any vector $x\in\{0,1\}^{N}$, only a constant $\varepsilon$ fraction of the constraints encoded by $A$ and $y$ are satisfied, $\Omega(N)$ queries are necessary.
For completeness, we will briefly describe the proof of the above lemma in Appendix B. The authors in [bogdanov2002lower] proved the following lower bound for testing E$(3,c)$LIN-2.
@@
For every $\alpha>0$, there are constants $c$ and $\delta>0$ such that every algorithm that distinguishes satisfiable instances of E$(3,c)$LIN-2 with $N$ variables from instances that are $(1/2-\alpha)$-far from satisfiable must have classical query complexity at least $\delta N$.
The key insight that is used to prove the above lemma is that applying Lemma 5.4, any $\delta N$ rows of $A$ are linearly independent, thus any subset of $\delta N$ entries of $Ay=z$ will look uniformly random. Hence $z$ is “$k$-wise independent” with $k=\delta N$. We will not formally prove the above lemma here, please refer to [bogdanov2002lower] for a formal proof.
Finally, we have the reduction that maps satisfying instances of testing E$(3,c)$LIN-2 to satisfying instances of testing $3$-colorability and vice-versa.
@@
There exists a reduction $\varphi$ that maps instances of testing E$(3,c)$LIN-2 to instances of testing 3-colorability such that the following hold:
@@
    If an input $x$ to E$(3,c)$LIN-2 is satisfiable, then $\varphi(x)$ is a $3$-colorable graph;
@@
    If an input $x$ to E$(3,c)$LIN-2 is far from being satisfiable, then $\varphi(x)$ is a graph that is far from being $3$-colorable.
5.2 Quantum lower bound for testing E$(3,c)$LIN-2 and $3$-colorability
We will first prove the quantum lower bound for testing E$(3,c)$LIN-2. Our result is stated as follows.
@@ 5.2 Quantum lower bound for testing E$(3,c)$LIN-2 and $3$-colorability
For every $\alpha>0$ there are constants $c$ and $\delta>0$ such that every algorithm that distinguishes satisfiable instances of E$(3,c)$LIN-2 with $N$ variables from instances that are $(1/2-\alpha)$-far from satisfiable must have quantum query complexity at least $\delta N/2$.
@@
Now let us prove Lemma 5.8.
Proof of Lemma 5.8.
Following Lemmas 5.4 and 5.5, we know that there exists a matrix $A$ whose $\delta N$ rows are linearly independent, for which testing E$(3,c)$LIN-2 requires $\Omega(N)$ classical queries. Moreover, from Proposition 5.9, we know that any quantum algorithm that performs less than $k/2$ queries can not distinguish a uniformly random vector from a $k$-wise independent vector. Now let us set $k=\delta N$. Combining all the above, this implies that at least $\delta N/2$ quantum queries are necessary for testing E$(3,c)$LIN-2. ∎
Now we are finally ready to prove Theorem 5.1.
Proof of Theorem 5.1.
From Lemma 5.8, we know that the quantum query complexity of testing E$(3,c)$LIN-2 is $\Omega(N)$. In order to prove similar lower bound for testing $3$-colorability, we will again use a reduction approach. Given a pair of hard instances corresponding to testing E$(3,c)$LIN-2, we will apply the reduction $\psi$ mentioned in Lemma 5.7. Similar to the classical setting, $\varphi$ will map the yes instances of E$(3,c)$LIN-2 to instances of $3$-colorable graphs and vice-versa. So, the quantum query lower bound of $\Omega(N)$ carries forward from E$(3,c)$LIN-2 to $3$-colorability. Thus, we conclude that $\Omega(N)$ quantum queries are necessary to test $3$-colorability in the bounded degree model. ∎

### v3

@@ 4.3 Main lower bound
Let us fix $f=(\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k})$ with domain $D=D_{(\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k})}$ (see Proposition˜4.12). For technical reasons, in the rest of the section we fix $k\geq 3$ and $N=\lceil 20(2k)^{k/2}\rceil R$. [^3] These parameters are used in [polynomial_strikes_back] to prove Proposition A.4, which we will use.
@@
This subsection is dedicated to the proof of the following lemma which, together with Lemma˜4.10, implies Theorem˜4.1. Observe that we have to zero out the support of the dual polynomial outside of $H_{\leq N}^{NR}$, since our target domain is not $D$ but $D\cap H_{\leq N}^{NR}$ in Lemma˜4.10.
@@
The construction of $\zeta$ starts by block composing (Definition˜4.13) two dual polynomials $\phi,\psi$, one for $\mathrm{GapOR}_{R}^{\gamma}$ and one for $\mathrm{THR}_{N}^{k}$. The dual polynomial $\phi$ for $\mathrm{GapOR}_{R}^{\gamma}$ is given by Proposition˜4.15. The dual polynomial $\psi$ for $\mathrm{THR}_{N}^{k}$ is given by the first part of Lemma˜4.16.
The block composition $\phi\star\psi$ is a good candidate for the dual polynomial of $f$. Indeed, Lemma˜4.18 shows that it satisfies Equation˜8, showing correlation at least $9/10>2/3$. One could also check that it satisfies Equation˜9. Nonetheless it does not satisfy Equation˜7.
We can now use the second part of Lemma˜4.16 to argue that there exists another dual polynomial $\zeta$ that satisfies Equation˜7 and Equation˜9. Moreover, this $\zeta$ is close to $\phi\star\psi$ so that it also satisfies Equation˜8, with the weaker but sufficient correlation 9/102/9>2/39/10-2/9>2/3. This concludes the proof. ∎
@@
The second lemma is the rephrasing of several scattered results in [polynomial_strikes_back] that we unify into one statement for more clarity. For the sake of completeness, we explain how to prove it in Section˜A.2.
@@
Let $N=\lceil 20(2k)^{k/2}\rceil R$ and $\phi:\{-1,1\}^{R}\to\mathbb{R}$ from Proposition˜4.15. Then there exist $\psi:\{-1,1\}^{N}\to\mathbb{R}$ and $\zeta:\{-1,1\}^{NR}\to\mathbb{R}$ such that
@@
For the next lemma, we will use the following proposition, which was implicitly used in the proofs of [polynomial_strikes_back, Propositions 5.5 and 5.6], but not stated in this general form. We include its proof in Section˜A.3. By convention, we denote $D_{+1}=D_{+}$ and $D_{-1}=D_{-}$.
@@
As now we focus on the composed function $f$ (and the dual composition $\phi\star\psi$), the domain is not restricted to small Hamming weight inputs anymore.
@@
Let $N=\lceil 20(2k)^{k/2}\rceil R$ and $0<\gamma<1/4^{k-1}$. Functions $\phi$ from Proposition˜4.15 and $\psi$ from Lemma˜4.16 satisfy
@@
Proof of claim. Using Item 2 of Proposition˜4.17, the left hand side can be written as
@@
Recall that $\phi(z)=0$ when $z$ is anything but $1^{R}$ or $1^{R}$, so only two terms are left to study.
If $z=-1^{R}$, using Item 1 of Lemma˜4.16 each $y_{i}$ is 1-1 with probability $\leq 1-1/4^{k-1}$ and 1 with probability $\geq 1/4^{k-1}$. If there is any $y_{i}=1$ then the value of the OR is still 1-1. The probability of this event is $\geq 1-(1-1/4^{k-1})^{R}\geq 1-e^{-\frac{R}{4^{k-1}}}$. So the expected value is $\leq(-1)(1-e^{-\frac{R}{4^{k-1}}})+e^{-\frac{R}{4^{k-1}}}=-1+2e^{-\frac{R}{4^{k-1}}}$. Since in this case $\phi(-1^{R})=-1/2$, the contribution to the sum is at most $1/2-e^{-\frac{R}{4^{k-1}}}$.
If $z=1^{R}$, then, using Item 1 of Lemma˜4.16 again, each $y_{i}$ is 1-1 with probability $\leq 1/(48N)$. If any $y_{i}$ is 1-1 then the value of the OR becomes 1-1. The union bound tells us that the probability of this is $\leq R/(48N)$, so the expected value is at least R/(48N)+1R/(48N)=1R/(24N)-R/(48N)+1-R/(48N)=1-R/(24N). Multiplied by $\phi(1^{R})=1/2$ the contribution is at least 1/2R/(48N)1/2-R/(48N). Thus, the first term can be lower bounded by $1-e^{-\frac{R}{4^{k-1}}}-\frac{R}{48N}$. $\diamond$
@@
Proof of claim. By Item 1 of Proposition˜4.17 with $S=\{-1,1\}^{NR}\setminus D$, the term can be written as follows,
@@
which, using that $|\phi(z)|=1/2$ when $z$ is $1^{R}$ or $1^{R}$ and 0 otherwise, collapses to
$\Pr_{x\sim\lambda^{\otimes R}}[x\notin D\mid(\dots,\mathrm{sgn}(\psi(x_{i})),z\dots)=-1^{R}]+\Pr_{x\sim\lambda^{\otimes R}}[x\notin D\mid(\dots,\mathrm{sgn}(\psi(x_{i})),\dots)=1^{R}].$
@@
Similarly, for the second term, if we sample $x_{i}$ from the conditional distribution $(\lambda|\mathrm{sgn}(\psi(x_{i}))=-1)$, then $\Pr[q_{i}=1]=\Pr[\mathrm{THR}_{N}^{k}(x_{i})=1|\mathrm{sgn}(\psi(x_{i}))=-1]=2\sum_{x_{i}\in D_{-}}|\psi(x_{i})|\leq 1-\frac{1}{4^{k-1}}$ (for the last step we used Item 1 of Lemma˜4.16 again).
Then, similarly to the first term, we can upper bound the probability. Now in the last step we use the Chernoff bound, which introduces the constraint $\gamma<\frac{1}{4^{k-1}}$.
@@
Finally, we can conclude the proof of Theorem˜4.1.
Proof of Theorem˜4.1.
By Lemma˜4.14, there is a dual polynomial for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ of pure high degree $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ that is only supported on $H_{\leq N}^{NR}$. By Theorem˜4.8 this means that the double-promise $\delta$-approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain restricted to $H_{\leq N}^{NR}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$. Using Lemma˜4.10 with $c=\lceil 20(2k)^{k/2}\rceil$, we obtain that the bounded-error quantum query complexity of $\mathrm{Collision}_{N,N}^{k,\gamma^{\prime}}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ if $\gamma^{\prime}=\gamma/c<1/(4^{k-1}\lceil 20(2k)^{k/2}\rceil)$. This implies the same lower bound on testing $k$-collision-freeness with $\varepsilon=\gamma^{\prime}$ as $\mathrm{Collision}$ is just a more restricted version of the same problem. ∎
@@ 5 Testing $3$-colorability
Theorem 5.1 (Restatement of Theorem˜1.4).
@@ 5.1 Classical lower bound approach for testing $3$-colorability
As we discussed in the overview, to prove the lower bound of $3$-colorability, the authors in [bogdanov2002lower] studied another problem called E(3,c)(3,c)LIN-2, a problem related to deciding the satisfiability of a system of linear equations. Then the authors designed a reduction to $3$-colorability from E(3,c)(3,c)LIN-2, which finally proves the linear query complexity lower bound for testing $3$-colorability. We will also follow a similar approach here. Let us first formally define the problem of E(3,c)(3,c)LIN-2, where below $\mathbb{F}_{2}$ denotes the field with two elements.
Definition 5.3 (E(3,c)(3,c)LIN-2).
@@
Using the existence of the matrix $A$ corresponding to Lemma˜5.4, the authors in [bogdanov2002lower] used Yao’s minimax lower bound technique to prove a linear lower bound for testing E(3,c)(3,c)LIN-2. More formally, they designed a pair of hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$, such that unless $\Omega(N)$ queries are performed, no algorithm can distinguish between them.
@@
There exists a matrix $A\in\{0,1\}^{cN\times N}$ (similar to the matrix mentioned in Lemma˜5.4) such that given a parameter $\varepsilon\in(0,1)$ and query access to $A$ and a vector $y\in\{0,1\}^{cN}$, in order to distinguish if there exists another vector $x\in\{0,1\}^{N}$ such that $Ax=y$, or for any vector $x\in\{0,1\}^{N}$, only a constant $\varepsilon$ fraction of the constraints encoded by $A$ and $y$ are satisfied, $\Omega(N)$ queries are necessary.
For completeness, we will briefly describe the proof of the above lemma in Appendix˜B. The authors in [bogdanov2002lower] proved the following lower bound for testing E(3,c)(3,c)LIN-2.
@@
For every $\alpha>0$, there are constants $c$ and $\delta>0$ such that every algorithm that distinguishes satisfiable instances of E(3,c)(3,c)LIN-2 with $N$ variables from instances that are $(1/2-\alpha)$-far from satisfiable must have classical query complexity at least $\delta N$.
The key insight that is used to prove the above lemma is that applying Lemma˜5.4, any $\delta N$ rows of $A$ are linearly independent, thus any subset of $\delta N$ entries of $Ay=z$ will look uniformly random. Hence $z$ is “$k$-wise independent” with $k=\delta N$. We will not formally prove the above lemma here, please refer to [bogdanov2002lower] for a formal proof.
Finally, we have the reduction that maps satisfying instances of testing E(3,c)(3,c)LIN-2 to satisfying instances of testing $3$-colorability and vice-versa.
@@
There exists a reduction $\varphi$ that maps instances of testing E(3,c)(3,c)LIN-2 to instances of testing 3-colorability such that the following hold:
@@
    If an input $x$ to E(3,c)(3,c)LIN-2 is satisfiable, then $\varphi(x)$ is a $3$-colorable graph;
@@
    If an input $x$ to E(3,c)(3,c)LIN-2 is far from being satisfiable, then $\varphi(x)$ is a graph that is far from being $3$-colorable.
5.2 Quantum lower bound for testing E(3,c)(3,c)LIN-2 and $3$-colorability
We will first prove the quantum lower bound for testing E(3,c)(3,c)LIN-2. Our result is stated as follows.
@@ 5.2 Quantum lower bound for testing E(3,c)(3,c)LIN-2 and $3$-colorability
For every $\alpha>0$ there are constants $c$ and $\delta>0$ such that every algorithm that distinguishes satisfiable instances of E(3,c)(3,c)LIN-2 with $N$ variables from instances that are $(1/2-\alpha)$-far from satisfiable must have quantum query complexity at least $\delta N/2$.
@@
Now let us prove Lemma˜5.8.
Proof of Lemma˜5.8.
Following Lemmas˜5.4 and 5.5, we know that there exists a matrix $A$ whose $\delta N$ rows are linearly independent, for which testing E(3,c)(3,c)LIN-2 requires $\Omega(N)$ classical queries. Moreover, from Proposition˜5.9, we know that any quantum algorithm that performs less than $k/2$ queries can not distinguish a uniformly random vector from a $k$-wise independent vector. Now let us set $k=\delta N$. Combining all the above, this implies that at least $\delta N/2$ quantum queries are necessary for testing E(3,c)(3,c)LIN-2. ∎
Now we are finally ready to prove Theorem˜5.1.
Proof of Theorem˜5.1.
From Lemma˜5.8, we know that the quantum query complexity of testing E(3,c)(3,c)LIN-2 is $\Omega(N)$. In order to prove similar lower bound for testing $3$-colorability, we will again use a reduction approach. Given a pair of hard instances corresponding to testing E(3,c)(3,c)LIN-2, we will apply the reduction $\psi$ mentioned in Lemma˜5.7. Similar to the classical setting, $\varphi$ will map the yes instances of E(3,c)(3,c)LIN-2 to instances of $3$-colorable graphs and vice-versa. So, the quantum query lower bound of $\Omega(N)$ carries forward from E(3,c)(3,c)LIN-2 to $3$-colorability. Thus, we conclude that $\Omega(N)$ quantum queries are necessary to test $3$-colorability in the bounded degree model. ∎